OpenStudy (anonymous):

22% of 50 is____ .a0 Help me plezz :(

OpenStudy (anonymous):

\[\frac{22}{100} \times 50 = ??\]

OpenStudy (anonymous):

22 % means 22/100 22/100 = x/50 isolate x = 22/100 * 50 x = 11

OpenStudy (anonymous):

\[22 \times \frac{50}{100} = 22 \times \frac{1}{2} = ??\]

OpenStudy (anonymous):

is it 11

OpenStudy (anonymous):

50 times 0.22 (22%) = 11

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

Yes it is 11.. Well Done..